Solve the equation. If necessary, round your answer to the nearest tenth.

4x^2 = 121

Answer by checkley79(3341)

4x^2 = 121
x^2=121/4
x^2=30.25
x=sqrt30.25
x=5.5 ans.
Proof:
4*5.5^2=121
4*30.25=121
121=121